# The physical basis of teleological evolutions

### v11

@@ 1.1 The physical basis of the teleological notion in the quantum domain
We should instead scale the quantum features down from the largest and longest-lasting evolution in existence, that of the universe. At the same time, of course, we should adopt quantum cosmology $\left[9\right]$, namely the assumption that the universe – a perfectly isolated system since by definition there is nothing outside it – evolves in a unitary quantum way.
@@
In fact, a consequent quantum interpretation of the Anthropic Principle $\left[10-13\right]$ would solve our problem. It would say that the evolution of the universe toward the “value” of life (”existence of sentient life” or “absence of life”) builds up a quantum correlation between the value (for values) of the fundamental physical constants since the beginning of time and the value of life today; this evolution could consequently occur with quantum speedup and would consequently be teleological in character.
@@
As is well known, the Fine-tuned Universe version of the Anthropic principle interprets the fact that the slightest change of the value of the fundamental constants since the beginning of time would have produced a radically different universe most probably unable to accommodate life.
@@
This version of the principle simply asserts that there is a correlation (we assume for simplicity one-to-one) between the value of the fundamental constants since the beginning of time and the value of life today. Of course, this correlation must have been built up by the evolution of the universe under the fundamental laws and constants. In fact, at the beginning of time, the value of the fundamental constants and that of life were unrelated: no matter the value of the fundamental constants, there could have been no life.
@@
The evolution of the (generic) universe toward the value of life can therefore occur with quantum speedup and consequently be teleological in character – the “can occur” is in the sense that there is the room for a quantum speedup.
In the universe we are in, what can occur with quantum speedup is of course the evolution of the universe toward life. The room for quantum speedup would then be inherited by the evolution of life on Earth, even if decohered. In fact, first, evolution times must be the same for the evolution of life on Earth, obtainable in principle from that of the universe by tracing out all the rest, and its image in the evolution of the universe. Second, the resources employed by the former evolution cannot exceed those employed by an image that comprises them.
In the end, being able to occur with quantum speedup, the evolution of life on Earth can be teleological; the same would consequently hold for the evolutions of species. Since it is reasonable, and conservative, to assume that the search for the value of life by the evolution of the universe under the fundamental physical laws and constants is random, its room for quantum speed should become an actual quantum speedup, as in quantum random search algorithms $\left[14\right]$.
This ends the second part of the work. Of course, we should keep in mind that, being under the quantum cosmological assumption, all the above is hypothetical.
Section 2. is about the teleological notion in the quantum framework – about the first part of the work and the Rationale; it can be seen as an extension and formalization of the corresponding part of the Rationale.
Section 3. is about the present, scientific, teleological notion in natural evolutions – the second part of the work and the Rationale.
@@ 2 The teleological notion in the quantum domain
This section provides the physical basis to the notion of teleological quantum evolution. It is in two parts. In the first we formalize the notion of completing the ordinary quantum description of quantum algorithm by time-symmetrizing it. This highlights the teleological character of quantum algorithms. The second is a generalization from quantum algorithms to quantum correlation.
@@ 2.1 Completing the ordinary quantum description of quantum algorithms by time-symmetrizing it
(2) This point is propaedeutic to point (4) below. We show how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary observer external to the quantum process. Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement: it would give her full information about the solution of the problem444 This is obvious in Grover algorithm, where the setting and the solution of the problem are both the number of the drawer with the ball. More in general, we must assume that the problem-solver does not know the problem-setting selected by the initial measurement but knows everything about the structure of the problem, thus also the relation between problem-setting and solution. .
@@ 2.1.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow555 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operation that would change nothing. .
@@
The ordinary quantum description of the quantum algorithm of Table I is with respect to the customary external observer, that is, an observer external to the quantum process. It is not a description for the problem-solver Alice. In fact, it would tell her, immediately after the initial measurement of $\hat{B}$, the problem-setting selected by the initial measurement (by the way, here they are both the number of the drawer with the ball). Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement.
@@ 2.1.3 Time-symmetrizing the ordinary quantum description of the quantum algorithm
Note that the selection, by the initial measurement, of half of the information that specifies the problem-setting is made possible by two facts: (i) the measurement occurs in a quantum superposition of all the possible problem settings, therefore it is possible to reduce it to the measurement of any partition of all the possible problem settings and (ii) the problem-setting never changes along the unitary evolution between the two measurements, therefore the measurement of the partition of problem-settings can be postponed along that evolution until making it simultaneous with the final measurement; this yields any desired sharing of the selection. Of course, as the even sharing of the selection can be performed in many possible mutually exclusive ways, eventually we should take their quantum superposition.
@@ 2.1.4 Relativizing the causal loops with respect to the problem-solver
The causal loops generated by time-symmetrizing the ordinary quantum description of the quantum algorithm, which is with respect to the external observer, do not yet map on those logically implied by the sheer existence of the quantum speedup, which is with respect to the problem-solver.
Therefore, we should relativize with respect to the problem-solver the causal loop of Table III by postponing after her problem-solving action the projection of the quantum state due to the initial measurement of $\hat{B}_{l}$ (whose outcome should be concealed from her). This yields the causal loop in the following table:
@@
The projection of the initial state of register $B$ due to the measurement of $\hat{B}_{l}$ – top-left corner of the diagram – is postponed after Alice’s problem-solving action, namely outside the present table that is limited to that action. With this postponement, the initial state of register $B$ goes unaltered through the initial measurement of $\hat{B}_{l}$ becoming, by $\hat{\mathbf{U}}_{1,2}$, the quantum superposition of tensor products above the vertical arrow. Alice’s final measurement of $\hat{A_{r}}$, which in the present time-symmetrization instance selects the value $1$ of the right bit of the number contained in register $A$, projects the state above the vertical arrow on the state below it. Then (in causal order), this final measurement outcome propagates backward in time by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until becoming the definitive outcome of the initial measurement. This backward in time propagation is an instance of the time-symmetrized description of the quantum algorithm with respect to the problem-solver.
@@
Also note that, for the mathematical equivalence between $\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$ and $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$ and disregarding the time-direction of causality, the bottom line of the diagram of Table IV can also be read in the usual left to right way:
@@
The entire causal loop is the zigzag diagram of table IV. It is a causal loop that changes its own past. The state of register $B$ in the top-left part of the diagram, which tells us that Alice is completely ignorant of the number of the drawer with the ball, changes into the state of register $B$ in the bottom-left part of the diagram, which tells us that Alice knows in advance, before beginning her problem-solving action, that the ball is in drawer either $01$ or $11$. Note that the bottom-line of the diagram represents the time-symmetrized quantum algorithm, mathematically equivalent to the diagram of Table V.
Since these causal loops obey classical logic, the number of drawer openings needed to find the solution in the optimal quantum case is that needed in the optimal logical way according to Table V diagram, namely just one. In other words, the computational complexity of the problem to be solved by Alice reduces. The problem of locating the ball in one of $4$ drawers – as for the state of register $B$ in the upper-left corner of the diagram of Table IV – becomes that of locating it in one of $2=\sqrt{4}$ drawers – as for the state of register $B$ in the bottom-left corner of Table IV diagram (or the left of Table V diagram). In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place quadratically reduces from $4$ basis vectors to $2=\sqrt{4}$.
@@ 2.2 Generalization
First, one can see that everything immediately generalizes to any number of drawers: it suffices to increase in an even way the number of qubits in each quantum register. By the way, since these qubits should be divided into two possible halves, each measured by each measurement, for simplicity we should assume that their number is even.
Note that we never needed to look inside $\hat{\mathbf{U}}_{1,2}$, we have just been dealing with the generic unitary transformation that builds up one-to-one correlation between two initially unrelated commuting observables without changing the eigenvalue of the observable selected by the initial measurement, see Table I. This, of course, covers any quantum algorithm where the notions of problem-setting and solution of the problem are well defined and the solution is determined with certainty. If the solution is not an invertible function of the problem-setting, it suffices that, at the end, the problem-solver, before reading the solution, measures the problem-setting. Of course, in any such quantum algorithm, the problem-solver always knows in advance half of the information about the solution. In $\left[6\right]$, we have seen that, when the problem is structured, this advanced knowledge can create the room for an exponential speedup.
The thing generalizes also beyond quantum algorithms. Let us consider a unitary evolution between an initial and a final one-to-one correlated observables that does not change the eigenvalue of the observable selected by the initial measurement. In this kind of process, one is free to postpone the initial measurement at will. When made simultaneous with the final measurement, the time-symmetrization we are dealing with must already be in place to become the perfect atemporal symmetry between the simultaneous measurements of two maximally entangled observables – see Table II. This is for the principle of sufficient reason: there would be no reason for the prevalence of either measurement in the simultaneous selection of the pair of one-to-one correlated measurement outcomes.
@@
(i) The unitary evolution builds up one-to-one correlation between two initially unrelated observables. This is, of course, the case of quantum algorithms but it can also be that of processes of identical formalization (as in Table I) that are not quantum algorithms. In any case the building up of the quantum correlation can occur with quantum speedup. If the process is not a quantum algorithm, the notion of advanced knowledge, on the part of the problem-solver, of half of the information about the solution she will produce and measure in the future might loose any meaning; however, it can be replaced by the notion of an equivalent reduction of the dimension of the Hilbert space in which the search for one-to-one correlation takes place.
@@
The notion that computation is possible in the quantum world and the fundamental notion of quantum bit was introduced by Finkelstein $[19]$ in 1969; that was in the context of seeing the universe as a quantum computer – indeed a quantum cellular automaton. In 1982, Feynman $\left[20\right]$ highlighted the higher than classical computational efficiency of the quantum world, noticing that the classical simulation of a quantum process can require an amount of physical resources exponentially higher than the amount involved in the quantum process itself. In 1985, Deutsch $\left[21\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup.
@@ 3 The possible physical basis of the teleological evolutions of life
Of course, a difficulty is seeing quantum features in evolutions of life that can have the size that the living system of Earth and have durations of billions of years. We have seen in the Rationale that the only possible way out is deriving these features from an assumed quantum evolution of the universe.
A consequent quantum interpretation of the Fine-tuned Universe version of the Anthropic Principle would say that the evolution of the universe toward the value of life (absence of life or existence of sentient life) under the fundamental physical laws and constants would build up a quantum correlation between the value of the fundamental constants since the beginning of time and the value of life today. The evolution in question could therefore occur with quantum speedup and would consequently be teleological in character. In the universe we are in, of course, the same would apply to the evolution of the universe toward life, then to the evolution of life on Earth, and eventually the evolutions of species. Like in the case of random quantum search algorithms $\left[14\right]$, under the (conservative) assumption that the search for one-to-one correlation (for the value of life) is random the room for quantum speedup would become an actual quantum speedup.
In the following subsections, after (trivially) encoding a toy model of the evolution of the universe toward the value of life on the quantum description of Grover quantum search algorithm, we examine the impact that a teleological character of the evolution of the living would have on different areas of science and knowledge.
@@ 3.1 Teleological evolution of the universe toward the value of life – a toy model
We should imagine a value of the fundamental constants observable $\hat{B}$ and a value of life observable $\hat{A}$. Say that the eigenvalue $01$ of $\hat{B}$ is the value of the fundamental constants compatible with the development of sentient life and the eigenvalues $00,10,11$ the values of those incompatible. We assume that, correspondingly, selecting the eigenvalue $01$ by the measurement of $\hat{A}$ at the present time would mean existence of sentient life; the eigenvalues $00,10,11$ would mean instead absence of life.
@@
Of course, this is exactly the quantum search algorithm described in the previous section, the version with respect to the problem-solver of Table II. As shown by its time-symmetrization (Table IV), it can undergo with a quantum speedup, and does even if the quantum search for the solution is random $\left[14\right]$.

### v12

@@ 1.1 The physical basis of the teleological notion in the quantum domain
This very enormity suggests another possibility. It is to adopt quantum cosmology $\left[9\right]$, namely the assumption that the universe – a perfectly isolated system since by definition there is nothing outside it – evolves in a unitary quantum way. We could then think of scaling the quantum features down from the largest and longest-lasting evolution in existence, that of the universe.
@@
In fact, a consequent quantum interpretation of the Anthropic Principle $\left[10-13\right]$ would provide a solution to our problem. It would say that the evolution of the universe toward the “value” of life (in particular, “existence of sentient life” or “absence of life”) builds up a quantum correlation between the value (for values) of the fundamental physical constants since the beginning of time and the value of life today; this evolution could consequently occur with quantum speedup and would thus be teleological in character.
@@
Let us see this in more detail. As is well known, the Fine-tuned Universe version of the Anthropic principle is a way of expressing the fact that the slightest change of the value of the fundamental constants since the beginning of time would have produced a radically different universe most probably unable to accommodate life.
@@
This version of the principle just asserts that there is a correlation (we assume for simplicity one-to-one) between the value of the fundamental constants since the beginning of time and the value of life today. Of course, this correlation must have been built up by the evolution of the universe under the fundamental laws and constants. In fact, at the beginning of time, the value of the fundamental constants and that of life were unrelated: no matter the value of the fundamental constants, there could have been no life.
@@
The evolution of the universe toward the value of life can therefore occur with quantum speedup and consequently be teleological in character – the “can occur” is said in the sense that there is certainly the room for its occurring with a quantum speedup444 Note that, in problem-solving, advanced knowledge on the part of the problem-solver of half of the information about the solution does not imply a quantum speedup but means room for it; in order that there is a quantum speedup, the quantum algorithm must benefit of this advanced knowledge. .
In the universe we are in, what can occur with quantum speedup is of course the evolution of the universe toward life. This room for quantum speedup is then inherited by the evolution of life on Earth, even if decohered. In fact, first, evolution times must be the same for the evolution of life on Earth, obtainable in principle from that of the universe by tracing out all the rest, and its image in the evolution of the universe. Second, the resources employed by the former evolution cannot exceed those employed by an image that comprises them.
In the end, being able to occur with quantum speedup, the evolution of life on Earth can be teleological; the same would consequently hold for the evolutions of life on Earth, in particular those of species. Since it is reasonable, and conservative, to assume that the search for the value of life by the evolution of the universe under the fundamental physical laws and constants is random, its room for quantum speed should become an actual quantum speedup, as in quantum random search algorithms $\left[14\right]$.
This ends the second part of the work. Of course, we should keep in mind that, being under the quantum cosmological assumption, this part is hypothetical.
Section 2. is about the teleological notion in the quantum framework; it can be seen as an extension and formalization of the first part of the Rationale.
Section 3. is about the present, scientific, teleological notion in natural evolutions – it is an extension of the second part of the Rationale.
@@ 2 The teleological notion in the quantum domain
This section provides the physical basis to the notion of teleological evolution in the quantum domain. It is in two parts. In the first, we formalize the notion of completing the ordinary quantum description of quantum algorithm by time-symmetrizing it. This is what highlights the teleological character of quantum algorithms. The second is the generalization from quantum algorithms to quantum correlation.
@@ 2.1 Completing the ordinary quantum description of quantum algorithms by time-symmetrizing it
(2) This point is propaedeutic to point (4) below. We show how to relativize with respect to the problem-solver the ordinary quantum description of the quantum algorithm, which is with respect to the customary observer external to the quantum process. Of course, the problem-solver must be concealed from the problem-setting selected by the initial measurement: it would give her full information about the solution of the problem555 This is obvious in Grover algorithm, where the setting and the solution of the problem are both the number of the drawer with the ball. More in general, we must assume that the problem-solver does not know the problem-setting selected by the initial measurement but knows everything about the structure of the problem, thus also the relation between problem-setting and solution. .
@@ 2.1.1 Extending the usual quantum description of quantum algorithms to the process of setting the problem
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random, say it comes out $01$ – see the state of register $B$ under the vertical arrow666 By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operation that would change nothing here. .
@@
The ordinary quantum description of the quantum algorithm of Table I is with respect to the customary external observer, that is, an observer external to the quantum process. It is not a description for the problem-solver Alice. In fact, it would tell her, immediately after the initial measurement of $\hat{B}$, the problem-setting selected by the initial measurement and thus the solution of the problem. By the way, here they are both the number of the drawer with the ball. Of course, in the quantum description of the quantum algorithm with respect to Alice, she must be concealed from the outcome of the initial measurement.
@@ 2.1.3 Time-symmetrizing the ordinary quantum description of the quantum algorithm
Note that the selection, by the initial measurement, of half of the information that specifies the problem-setting is made possible by two facts: (i) the measurement occurs in a quantum superposition of all the possible problem settings, therefore it is possible to reduce it to the measurement of any partition of all the possible problem settings and (ii) the problem-setting never changes along the unitary evolution between the two measurement outcomes, therefore the measurement of the partition of problem-settings can be postponed along that evolution until making it simultaneous with the final measurement. Of course, as the even sharing of the selection can be performed in many possible mutually exclusive ways, eventually we should take their quantum superposition.
@@ 2.1.4 Relativizing the causal loops with respect to the problem-solver
The causal loops generated by time-symmetrizing the ordinary quantum description of the quantum algorithm, which is with respect to the external observer, do not yet map on those logically implied by the sheer existence of the quantum speedup, which are with respect to the problem-solver. We should then relativize the cuasal loops of Table III with respect to the problem-solver. We should just conceal from her the part of the problem-setting selected by the initial measurement. This is done by postponing after her problem-solving action the projection of the quantum state due to the initial measurement of $\hat{B}_{l}$ This yields the causal loop in the following table:
@@
The projection of the initial state of register $B$ due to the measurement of $\hat{B}_{l}$ – top-left corner of the diagram – is postponed after Alice’s problem-solving action, namely outside the present table which is limited to that action. With this postponement, the initial state of register $B$ goes unaltered through the initial measurement of $\hat{B}_{l}$ becoming, by $\hat{\mathbf{U}}_{1,2}$, the quantum superposition of tensor products above the vertical arrow. Alice’s final measurement of $\hat{A_{r}}$, which in this time-symmetrization instance selects the value $1$ of the right bit of register $A$, projects the state above the vertical arrow on the state below it. Then (in causal order), this final measurement outcome propagates backward in time by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until becoming the definitive outcome of the initial measurement. This backward in time propagation is an instance of the time-symmetrized description of the quantum algorithm with respect to the problem-solver. It tells her, before she begins her problem-solving action, one of the possible halves of the information about the solution.
@@
Also note that, for the mathematical equivalence between $\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$ and $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$, the bottom line of the diagram of Table IV can also be read in the usual left to right way:
@@
Summing up, the entire causal loop is the zigzag diagram of table IV. We can see that it is one of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. The state of register $B$ in the top-left part of the diagram tells us that Alice, before beginning her problem-solving action, is completely ignorant of the number of the drawer with the ball. Through the zig-zag, this state changes into the state of register $B$ in the bottom-left part of the diagram, which tells us that Alice, before beginning her problem-solving action, knows in advance that the ball is in drawer either $01$ or $11$. She can then solve the problem in an optimal logical way by opening just either drawer. In equivalent terms, the computational complexity of the problem to be solved by Alice quadratically reduces from the problem of locating the ball in one of $4$ drawers to that of locating it in $2=\sqrt{4}$ drawers. In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place quadratically reduces from $4$ basis vectors to $2=\sqrt{4}$ basis vectors.
@@ 2.2 Generalization
First, one can see that everything immediately generalizes to any number of drawers: it suffices to increase in an even way the number of quantum bits in each quantum register.
The second observation is that we never needed to look inside $\hat{\mathbf{U}}_{1,2}$. This unitary transformation is just defined by the fact that it must build up one-to-one correlation between two initially unrelated commuting observables without changing the eigenvalue of the observable selected by the initial measurement, see Table I. Besides Grover algorithm, we could be dealing with any quantum algorithm where the solution is an invertible function of the problem-setting and is determined with certainty. By the way, if the solution is not an invertible function of the problem-setting, it suffices that, at the end, the problem-solver, before reading the solution, measures the problem-setting. In any such quantum algorithm, the problem-solver always knows in advance half of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution with fewer computation steps. In the optimal quantum case, the number of computation steps is that needed to solve the problem in an optimal logical way benefiting of the advanced knowledge in question. Of course, this explains the quadratic quantum speedup of Grover quantum search algorithm, which solves an unstructured problem. In $\left[7\right]$, we have seen that, when the problem is structured, the advanced knowledge of half of the information about the solution can create the room for an exponential speedup.
Things generalize also beyond quantum algorithms. Let us consider a unitary evolution between an initial and a final one-to-one correlated observables that does not change the eigenvalue of the observable selected by the initial measurement. In this kind of process, one is free to postpone the initial measurement at will. When made simultaneous with the final measurement, the time-symmetrization we are dealing with must already be in place to become the perfect atemporal symmetry between the simultaneous measurements of two maximally entangled observables – see the states above and below the vertical arrow in Table II. This is for the principle of sufficient reason: there would be no reason for the prevalence of either measurement in the simultaneous selection of the pair of one-to-one correlated measurement outcomes.
@@
(i) The unitary evolution builds up one-to-one correlation between two initially unrelated observables. This is, of course, the case of quantum algorithms but it can also be that of processes of identical form (as that Table I but for a higher number of quantum bits) that are not quantum algorithms. In any case the building up of the quantum correlation can occur with quantum speedup. If the process is not a quantum algorithm, the notion of advanced knowledge, on the part of the problem-solver, of half of the information about the solution she will produce and measure in the future might loose any meaning. However, it can be replaced by the notion of an equivalent reduction of the dimension of the Hilbert space in which the search for one-to-one correlation takes place.
@@
The notion that computation is possible in the quantum world and the fundamental notion of quantum bit was introduced by Finkelstein $[19]$ in 1969; that was in the search for a unification between quantum mechanics and relativity theory, more specifically in the context of seeing the universe as a quantum computer. In 1982, Feynman $\left[20\right]$ highlighted the higher than classical computational efficiency of the quantum world, noticing that the classical simulation of a quantum process can require an amount of physical resources exponentially higher than the amount involved in the quantum process itself. In 1985, Deutsch $\left[21\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup.
@@ 3 The possible physical basis of the teleological evolutions of life
Of course, a difficulty is seeing quantum features in evolutions of life that can have the size of the living system of Earth and have durations of billions of years. We have seen in the Rationale that the only possible way out is deriving these features from an assumed quantum evolution of the universe.
A consequent quantum interpretation of the Fine-tuned Universe version of the Anthropic Principle would say that the evolution of the universe toward the “value” of life (in particular, “existence of sentient life” or “absence of life”) under the fundamental physical laws and constants would build up a quantum correlation between the value of the fundamental constants since the beginning of time and the value of life today. The evolution in question could therefore occur with quantum speedup, being consequently teleological in character. In the universe we are in, of course, the same would apply to the evolution of the universe toward life and then – even if decohered – to the evolution of life on Earth and consequently to the evolutions of species. Like in the case of random quantum search algorithms $\left[14\right]$, under the (conservative) assumption that the search for one-to-one correlation (for the value of life) is random, the room for quantum speedup would become an actual quantum speedup.
In the following subsections, after encoding a toy model of the evolution of the universe toward the value of life on the quantum description of Grover quantum search algorithm, we examine the impact that a teleological character of the evolution of the living would have on different areas of science and knowledge.
@@ 3.1 Teleological evolution of the universe toward the value of life – a toy model
We should imagine a value of the fundamental constants observable $\hat{B}$ and a value of life observable $\hat{A}$. Say that the eigenvalue $01$ of $\hat{B}$ is the value of the fundamental constants compatible with the development of sentient life and the eigenvalues $00,10,11$ the values of those incompatible. We assume that, correspondingly, selecting the eigenvalue $01$ by the measurement of $\hat{A}$ at the present time would mean existence of sentient life; selecting the eigenvalues $00,10,11$ would mean instead absence of life.
@@
Of course, this is exactly the quantum search algorithm described in the previous section, the version with respect to the problem-solver of Table II. As shown by its time-symmetrization (Table IV), it can undergo with a quantum speedup and does so if the quantum search for the solution is random $\left[14\right]$.